A penny is .22 m from the center horizontal turntable that rotates once every 1.9 s. If the coefficients of static and kinetic friction between the disk and the penny are .28 and .21, respectively, then will the penny slide. Show work and document the answer quantitatively.  
Answer #1

Centripetal acceleration

a = r ( 2 pi /t) ^2 = 2.403 m/s^2

Frictional acceleration

af = u g = 0.28* 9.8 = 2.744 m/s^2

Since, max static Frictional acceleration is greater than centripetal acceleration, so the object won't move.

=======

Comment in before rate in case any doubt, will reply for sure.. Goodluck u